OpenStudy (anonymous):

A rectangular prism has a length of 12 cm and a width of 4 cm. If the volume of the prism is 240 cm3, what is the height?

OpenStudy (anonymous):

justin beiber is gayyyyy

OpenStudy (anonymous):

vol = height * area 240 = h * 48 h = 240 / 48

OpenStudy (anonymous):

volume= length*breadth*height 240=12*4*h h=240/(12*4) h=5cm